Determine the roots of the function f(x) = (5^2x −6)^2 −(5^2x −6)−12.
1
Expert's answer
2020-11-17T05:41:20-0500

Let's note first "t:=5^{2x}-6" , then we can write:

"t^2-t-12=0"

It is an equation of second order so we can the solutions pretty easily (either with Vieta's formulas, either with the explicit formula "\frac{-b\pm\sqrt{b^2-4ac}}{2a}" ):

"t_1=4,t_2=-3"

Now let's come back from our substitution :

"5^{2x_1}-6=4, 5^{2x_2}-6=-3"

Therefore we find two solutions :

"x_1=\frac{1}{2}\log_5(10), x_2=\frac{1}{2}\log_5(3)"